Find the area bound by the curves y=-x^2+4 and y=0.
1
Expert's answer
2020-12-21T17:33:46-0500


Solve


"-x^2+4=0"

"x_1=-2, x_2=2"

We see that "y=-x^2+4\geq0" for "-2\leq x\leq2." Then


"Area=\displaystyle\int_{-2}^2(-x^2+4)dx=\big[-\dfrac{x^3}{3}+4x\big]\begin{matrix}
 2 \\
 -2
\end{matrix}"

"=-\dfrac{8}{3}+8-(\dfrac{8}{3}-8)=\dfrac{32}{3}(units^2)"

"\dfrac{32}{3}" square units